An uncharged capacitor and a resistor are connected in series to a source of emf. If e m f = 10.0 V, C = 21.0 µF, and R = 100 Ω, find the following.

(a) the time constant of the circuit
(b) the maximum charge on the capacitor
(c) the charge on the capacitor at a time equal to one time constant after the battery is connected
Answer #1

part a )

time constant = t

t = RC

t = 100 ohm * 21 x 10^-6 F = 2.1 x 10^-3 s = 2.1 ms

part b )

Qmax = C*V

V = 10 V

Qmax = 2.1 x 10^-4 C = 210 uC

part c )

Q = Qmax * (1-e^-T/t)

T = t

Q = Qmax(1-e^-1)

Q = 132.75 uF